OpenStudy (sloppycanada):

Five rational function equations to help me with, if you'd be oh so kind. (5x+1)/x < 1

OpenStudy (sloppycanada):



OpenStudy (anonymous):

Ok i have an idea but i might be wrong so hopefully this might help. https://www.khanacademy.org/math/algebra2/rational-expressions/solving-rational-equations/v/solving-rational-equations-2

OpenStudy (anonymous):

Im sorry ; ~;

OpenStudy (unklerhaukus):

\[\frac{5x+1}{x}<1\\ \frac{5x}x+\frac1x<1\\ \quad\dots\quad<1\]

OpenStudy (unklerhaukus):

the first term simplifies . . .

OpenStudy (sloppycanada):

5x^2 + 1/x <1?

OpenStudy (unklerhaukus):

not quite, the x's in the first term dont multiply, they cancel

OpenStudy (sloppycanada):

Oh. Okay so 5 + 1/x <1?

OpenStudy (unklerhaukus):

good, now the next step is to take away 5 form both sides

OpenStudy (unklerhaukus):

*from

OpenStudy (sloppycanada):

1/x < -4

OpenStudy (unklerhaukus):

yes

OpenStudy (unklerhaukus):

now if we multiply both sides by x . . .

OpenStudy (sloppycanada):

1 = -4x So x = -1/4

OpenStudy (unklerhaukus):

we get -1/4 < x but this is because we assumed that x was not negative (we didn't swap the direction of inequality)

OpenStudy (unklerhaukus):

if we had assumed that x was negative, we would have switched the direction of inequality, when we multiplied both sides by x, we would have gotten -1/4 > x

OpenStudy (unklerhaukus):

hmmm,

OpenStudy (sloppycanada):

Okay thanks. For this next one (assuming we're good here), wouldn't I just want to find the quadratic parts?

OpenStudy (sloppycanada):